# From Scores to Queues: Operationalizing Cross-Chain Obfuscation Signals for Smart-Contract Audits

### v1

@@ III-J Error Analysis
IV Cross-Chain Obfuscation Measurements and Findings
@@ IV-A Data and Measurement Setup
We also test direct transfer of the Ethereum cutoff 18.07 and compute on other chains (Table V).
@@ IV-B1 Cross-chain distributions and quantiles
Table IV reports score quantiles. Medians are similar (around 5.0–5.4), but high quantiles diverge: BSC p90/p95/p99 = 9.60/12.45/16.82, notably lower than Polygon (13.50/15.66/18.72) and Avalanche (14.09/16.30/19.18). This indicates a lighter tail on BSC and heavier tails on Polygon/Avalanche.
@@
Table V reports tail shares under the Ethereum cutoff (18.07): BSC 0.48%, Polygon 1.58%, and Avalanche 2.32%. Relative to Ethereum’s 1.00% baseline, BSC shrinks while Polygon/Avalanche inflate, showing that a fixed threshold does not translate consistently across chains.
@@
Practical reminder. Known incident contracts do not necessarily fall into the Top1% tail (see RQ4 percentiles). In monitoring scenarios, we therefore suggest maintaining a wider “watch band” (e.g., p95–p99) for continuous observation, while using the main/extreme cutoffs for resource-constrained prioritization.
@@
IV-B3 Threshold-transfer sensitivity across chain-specific cutoffs
@@ IV-C RQ2: Tail Structure
Following RQ1. If distributions differ, we ask what drives those differences: what actually dominates the high-obfuscation tail—standard ERC/NFT contracts, proxy templates, or low-visibility bespoke logic? We combine ERC labels, a proxy indicator, and signature-density evidence to build an interpretable profile.
@@ IV-C1 ERC20/ERC721 presence in the tail
Table VI compares overall vs. main-cutoff tail ERC shares. Overall ERC20 shares are 2.88% (BSC), 0.35% (Polygon), and 0.84% (Avalanche), while overall ERC721 shares are 0.07%, 3.42%, and 0.55%, respectively. In the tail, ERC20/ERC721 shares drop sharply: BSC ERC20 falls from 2.88% to 0.17 (17), and Polygon/Avalanche ERC721 drops to (order-of-magnitude decline).
@@
IV-C2 Proxy (heuristic indicator) prevalence
Scope note. We distinguish two notions: (i) EIP-1167 minimal proxy (strict template matching, extremely rare overall and in the tail, serving as a lower bound), and (ii) a broader proxy indicator (heuristic patterns covering transparent/UUPS/custom forwarders, potentially over/under-counting). Table VII reports (ii); strict minimal proxy is only a lower-bound reference and should not be interpreted as equivalent to EIP-1167 prevalence. We mark contracts with a proxy indicator and report prevalence in overall and tail sets (Table VII). The indicator is based on heuristic proxy bytecode patterns and forwarding behaviors, and is intentionally broad rather than a strict EIP-1167 detector; as such it may over/under-count. Proxy indicators remain rare overall (sub-1%); under the main cutoff, BSC shows clear enrichment (0.25%1.03%) while Polygon/Avalanche are flat or slightly lower. This suggests proxy patterns are not the dominant tail component, but can be active in certain chains’ high-score regions. Therefore, ERC/Proxy alone still cannot explain the tail, motivating structural cues below.
@@ IV-C4 Manual inspection of top tail exemplars
Consistency with high-risk audit cues. The above signals match common “high explanation-cost” audit cues: rare selectors and low signature density imply sparse, opaque interfaces; enriched external-call and return-data handling suggest more complex inter-contract orchestration; and cross-chain template reuse indicates conditions for bulk deployment and rapid propagation. We therefore treat the high-obfuscation tail as a higher-priority audit queue.
Finding 2 (RQ2). The high-score tail can be translated into an actionable audit profile: rare selectors + external-call enrichment + low signature density jointly point to hidden-logic, template-driven families. These structural signals can be used directly to triage the RQ1 queue and prioritize likely fast-spreading high-risk templates.
@@ IV-D1 Overall reuse and directional asymmetry
We use deduplicated runtime bytecode hashes as reuse identifiers and measure cross-chain overlap via Jaccard and overlap coefficients (Table VIII). Jaccard values are low (0.003–0.013), indicating that each chain has a largely distinct contract ecosystem. Figure 2 contrasts overall vs. tail reuse; “tail Jaccard” is computed after extracting tail sets (Top1%) within each chain and then computing hash Jaccard across tail sets, while “overall Jaccard” is computed on full deduplicated sets. Tail Jaccard is clearly higher than overall (roughly 0.0055–0.0188 vs 0.0034–0.0127), a 1.5–2.0 enrichment, with Polygon–Avalanche most pronounced; in some pairs it approaches a 2 increase (Fig. 2). Directionally, overlap is much higher from smalllarge chains than the reverse (e.g., AvalancheBSC 8.55% vs 0.36%, PolygonBSC 3.33% vs 0.42%), where and is the deduplicated hash set of chain . This asymmetry suggests cross-chain reuse is closer to diffusion from smaller ecosystems to larger ones (e.g., copying templates into larger markets for greater visibility and interaction opportunities). Together with the RQ2 profile, this supports the mechanism that template-like, low-visibility contracts are more reusable and thus amplified in cross-chain tails.
@@
IV-D2 High-obfuscation templates reuse across chains
@@
IV-D3 Cross-chain reuse cases
We observe identical bytecode hashes reappearing across multiple chains, and some cases even show identical addresses across chains. These patterns align with template-driven deployment or deterministic address strategies (e.g., CREATE2), providing traceable evidence of cross-chain reuse. Identical-address reuse further suggests deterministic deployment or unified scripting; defensively, this enables cross-chain tracing and preemptive alerting once a high-risk template is identified on any single chain.
@@ IV-E RQ4: Incident Alignment and Investigative Value
Following RQ3. Since high-score templates are reused across chains, RQ4 asks one practical question: do real cross-chain incidents fall into our high-score screening queue? If yes, the queue has real-world hit value; if no, it is only a weak clue. We assess this via incident alignment and case-based inspection.
@@ IV-E1 Alignment overview
We extract incident-related addresses from public reports and tx/log evidence and match them against deduplicated contracts. Coverage varies by chain (about 31% on BSC, lower on Polygon/Avalanche); tx/log-resolved evidence dominates. While alignment is incomplete, it is sufficient to test whether high-score queues hit real incidents.
@@ IV-E2 Percentile localization: real incidents hit the high-score queue
Table IX reports aligned samples and their within-chain percentiles (cross-chain spillover cases). During screening, we identified two cross-chain incident samples that fall into the p99 tail but not the p99.9 extreme tail. The picture is consistent: p99 acts as the practical main audit queue, while p99.9 is best reserved for the most urgent, minimal queue. In practice, should be used for prioritization: review p99 tail first and combine with reuse clusters and low-signature evidence for deeper triage.
@@ IV-E3 Case studies of cross-chain spillovers (excerpt)
Finding 4 (RQ4). The aligned cross-chain incident samples fall into the p99 tail but not p99.9, indicating that the high-score queue has real-world hit value and that p99 is the practical main audit queue. Transit Swap and New Free DAO provide cross-chain spillover evidence in public analyses, showing that high-obfuscation templates can leak across chains in practice. is therefore a usable prioritization signal to narrow down large contract populations, while p99.9 is best reserved for the most urgent, minimal queue.
@@ IV-F Summary
V Discussion
V-A Operational Implications and Audit Workflow
@@
V-B Mechanistic Interpretation of Drift and Spillover
@@
V-C Scope and Risk Boundaries
@@
V-D Limitations and Future Work
@@
VI Conclusion
We present HObfNET, a fast surrogate of Obfs_Tool whose inference cost and speed support million-scale, cross-chain scoring. We analyze four chains at scale (ETH 1,042,923; BSC 2,308,899; Polygon 288,611; Avalanche 96,173) and provide operational, within-chain thresholds (p99/p99.9; e.g., ETH 18.07/22.69, BSC 16.82/19.74, Polygon 18.72/20.51, Avalanche 19.18/20.67). On top of Ethereum supervision, we answer four research questions: (i) scores drift across chains; (ii) the high-score tail exhibits a structural profile with rare selectors, external-call enrichment, and low signature density, supporting secondary triage; (iii) high-score templates show directional cross-chain reuse, with tail Jaccard about 1.5–2.0 higher than overall; and (iv) incident alignment on cases (all in p99, none in p99.9) shows real-world hit value of the high-score queue within the aligned evidence set. Collectively, these results yield an actionable cross-chain audit workflow and prioritization queues for security operations.

### v2

@@ III-J Error Analysis
IV Results
@@ IV-A Data and Measurement Setup
Cross-chain tool alignment (10k per chain). To directly validate score comparability, we run ObfProbe on 10k contracts per chain and align tool scores with model predictions (Table IV). The agreement is strong (PCC 0.96–0.97, MAE 0.45–0.47), supporting as a comparable cross-chain signal for distribution and tail analysis.
| Chain | N | MAE | RMSE | MAPE | PCC |
| BSC | 10,000 | 0.45 | 0.93 | 5.7% | 0.97 |
| Polygon | 10,000 | 0.47 | 0.95 | 5.9% | 0.96 |
| Avalanche | 10,000 | 0.45 | 0.91 | 5.3% | 0.97 |
Feature definitions and heuristics. We extract 4-byte selectors from PUSH4 immediates in runtime bytecode; the number of unique selectors is the signature count. Signature density is defined as . ERC20/721 labels are assigned when bytecode contains a core selector set (ERC20: totalSupply, balanceOf, transfer, approve, transferFrom; ERC721: balanceOf, ownerOf, approve, setApprovalForAll, transferFrom, safeTransferFrom), used as a heuristic indicator rather than a definitive classification. The proxy indicator marks contracts whose runtime bytecode exhibits delegatecall-based forwarding patterns (e.g., presence of DELEGATECALL with short forwarding dispatch), and is intentionally broader than strict EIP-1167 matching. For selector/opcode enrichment, we compute with a small smoothing term in the denominator, and we only report patterns above a minimum occurrence threshold to avoid unstable lift ratios.
@@ IV-B1 Cross-chain distributions and quantiles
Table V reports score quantiles. Medians are similar (around 5.0–5.4), but high quantiles diverge: BSC p90/p95/p99 = 9.60/12.45/16.82, notably lower than Polygon (13.50/15.66/18.72) and Avalanche (14.09/16.30/19.18). This indicates a lighter tail on BSC and heavier tails on Polygon/Avalanche.
@@
We test direct transfer of the Ethereum cutoff 18.07 and compute on other chains; Table VI reports the tail shares: BSC 0.48%, Polygon 1.58%, and Avalanche 2.32%. Relative to Ethereum’s 1.00% baseline, BSC shrinks while Polygon/Avalanche inflate, showing that a fixed threshold does not translate consistently across chains.
@@
Practical reminder. Cross-chain spillovers are rare and may not always land in the extreme tail. We therefore treat p99 as the main queue and suggest maintaining a wider “watch band” (e.g., p95–p99) for continuous observation, while using the main/extreme cutoffs for resource-constrained prioritization.
@@
IV-B3 Threshold Transfer Sensitivity Across Chain-Specific Cutoffs
@@ IV-C RQ2: Tail Structure
Following RQ1. If distributions differ, we ask what drives those differences. Our goal is not just description: the tail profile should directly signal high risk. In practice, the high-obfuscation tail is dominated by low-visibility, hard-to-audit logic rather than standard ERC/NFTs or simple proxies. We therefore use ERC labels, a proxy indicator, and signature density to connect structure to risk: rare selectors + external-call enrichment + low signature density indicate hidden logic, cross-contract orchestration, and high audit cost.
@@ IV-C1 ERC20/ERC721 presence in the tail
Table VII compares overall vs. main-cutoff tail ERC shares. Overall ERC20 shares are 2.88% (BSC), 0.35% (Polygon), and 0.84% (Avalanche), while overall ERC721 shares are 0.07%, 3.42%, and 0.55%, respectively. In the tail, ERC20/ERC721 shares drop sharply: BSC ERC20 falls from 2.88% to 0.17% (17), and Polygon/Avalanche ERC721 drops to (order-of-magnitude decline).
@@
IV-C2 Proxy Indicator Prevalence
Scope note. We distinguish two notions: (i) EIP-1167 minimal proxy (strict template matching, extremely rare overall and in the tail, serving as a lower bound), and (ii) a broader proxy indicator (heuristic patterns covering transparent/UUPS/custom forwarders, potentially over/under-counting). Table VIII reports (ii); strict minimal proxy is only a lower-bound reference and should not be interpreted as equivalent to EIP-1167 prevalence. We mark contracts with a proxy indicator and report prevalence in overall and tail sets (Table VIII). The indicator is based on heuristic proxy bytecode patterns and forwarding behaviors, and is intentionally broad rather than a strict EIP-1167 detector; as such it may over/under-count. Proxy indicators remain rare overall (sub-1%); under the main cutoff, BSC shows clear enrichment (0.25%1.03%) while Polygon/Avalanche are flat or slightly lower. This suggests proxy patterns are not the dominant tail component, but can be active in certain chains’ high-score regions. Therefore, ERC/Proxy alone still cannot explain the tail, motivating structural cues below.
@@ IV-C4 Manual inspection of top tail exemplars
Consistency with high-risk audit cues. The above signals match common “high explanation-cost” audit cues and point to risk: rare selectors and low signature density imply sparse, opaque interfaces; enriched external-call and return-data handling suggest complex inter-contract orchestration; and cross-chain template reuse enables fast propagation. In other words, the tail profile is not just structure—it is a high-risk signature. We therefore treat the high-obfuscation tail as a higher-priority audit queue.
Finding 2 (RQ2). The high-score tail is a high-risk signature, not just a structural description. Rare selectors + external-call enrichment + low signature density jointly indicate hidden logic, cross-contract manipulation, and high audit cost—exactly the traits we associate with risky contracts. These signals can be used directly to triage the RQ1 queue and prioritize likely fast-spreading high-risk templates.
@@ IV-D1 Overall Reuse and Directional Asymmetry
We use deduplicated runtime bytecode hashes as reuse identifiers and measure cross-chain overlap via Jaccard and overlap coefficients (Table IX). Jaccard values are low (0.003–0.013), indicating that each chain has a largely distinct contract ecosystem. Figure 3 contrasts overall vs. tail reuse; “tail Jaccard” is computed after extracting tail sets (Top1%) within each chain and then computing hash Jaccard across tail sets, while “overall Jaccard” is computed on full deduplicated sets. Tail Jaccard is clearly higher than overall (roughly 0.0055–0.0188 vs 0.0034–0.0127), a 1.5–2.0 enrichment, with Polygon–Avalanche most pronounced; in some pairs it approaches a 2 increase (Fig. 3). Directionally, overlap is much higher from smalllarge chains than the reverse (e.g., AvalancheBSC 8.55% vs 0.36%, PolygonBSC 3.33% vs 0.42%), where and is the deduplicated hash set of chain (Fig. 4). This asymmetry suggests cross-chain reuse is closer to diffusion from smaller ecosystems to larger ones (e.g., copying templates into larger markets for greater visibility and interaction opportunities). Together with the RQ2 profile, this supports the mechanism that template-like, low-visibility contracts are more reusable and thus amplified in cross-chain tails.
@@
IV-D2 High-Obfuscation Template Reuse Across Chains
@@
Cross-chain reuse cases. We observe identical bytecode hashes reappearing across multiple chains, and some cases even show identical addresses across chains. These patterns align with template-driven deployment or deterministic address strategies (e.g., CREATE2), providing traceable evidence of cross-chain reuse. Identical-address reuse further suggests deterministic deployment or unified scripting; defensively, this enables cross-chain tracing and preemptive alerting once a high-risk template is identified on any single chain.
@@ IV-E RQ4: Incident Alignment and Investigative Value
Following RQ3. RQ4 asks one practical question: for the alignable cross-chain spillover cases, does the high-score queue hit real risk? Public cross-chain spillovers are rare, so we use the two alignable cases as a lower-bound validation. If they hit the queue, it indicates real-world value; if not, the queue is only a weak clue.
@@ IV-E1 Alignment overview
We extract incident-related addresses from public reports and tx/log evidence and match them against deduplicated contracts. Public cross-chain spillovers are rare; the two alignable cases provide a lower-bound check of whether the high-score queue hits real incidents.
@@ IV-E2 Percentile localization: real incidents hit the high-score queue
Table X reports aligned samples and their within-chain percentiles (cross-chain spillover cases). During screening, we identified two cross-chain incident samples that fall into the p99 tail but not the p99.9 extreme tail. The picture is consistent: p99 acts as the practical main audit queue, while p99.9 is best reserved for the most urgent, minimal queue. In practice, should be used for prioritization: review p99 tail first and combine with reuse clusters and low-signature evidence for deeper triage.
@@ IV-E3 Case studies of cross-chain spillovers (excerpt)
Finding 4 (RQ4). The aligned cross-chain spillover samples fall into the p99 tail but not p99.9, indicating that the high-score queue has real-world hit value and that p99 is the practical main audit queue. Together with the smalllarge diffusion in RQ3, this implies that high-obfuscation templates can spill over and lead to costly incidents. In practice, use p99 as the main queue, keep p99.9 for emergencies, and trigger cross-chain tracing once a high-obfuscation reuse cluster appears on smaller chains.
@@ IV-F Summary
V Related Work
Bytecode analysis and decompilation. Large-scale bytecode analysis is enabled by advances in decompilation and function recovery. Gigahorse introduced a declarative decompiler that lifts EVM bytecode to higher-level Intermediate Representation and supports scalable analyses [9]. Elipmoc improves recovery for complex control flow and optimized bytecode, while Shrnkr further increases scalability with shrinking context sensitivity [10, 17]. Neural-FEBI improves function identification directly on EVM bytecode, enabling more precise downstream analyses [11, 8, 27, 14, 6]. These tools make obfuscation measurement feasible at scale and enable reuse analysis, but they are not designed to study cross-chain drift and have not been systematically applied to cross-chain distribution and reuse measurement.
VI Discussion
VI-A Operational Implications and Audit Workflow
@@
VI-B Mechanistic Interpretation of Drift and Spillover
@@
VI-C Scope and Risk Boundaries
@@
VI-D Limitations and Future Work
@@
VII Conclusion
We present HObfNET, a fast surrogate of Obfs_Tool whose inference cost and speed support million-scale, cross-chain scoring. We analyze four chains at scale (ETH 1,042,923; BSC 2,308,899; Polygon 288,611; Avalanche 96,173) and provide operational, within-chain thresholds (p99/p99.9; e.g., ETH 18.07/22.69, BSC 16.82/19.74, Polygon 18.72/20.51, Avalanche 19.18/20.67). On top of Ethereum supervision, we answer four research questions: (i) scores drift across chains; (ii) the high-score tail exhibits a structural profile with rare selectors, external-call enrichment, and low signature density, supporting secondary triage; (iii) high-score templates show directional cross-chain reuse, with tail Jaccard about 1.5–2.0 higher than overall; and (iv) incident alignment on the available cross-chain cases (all in p99, none in p99.9) shows real-world hit value of the high-score queue within the aligned evidence set. Collectively, these results yield an actionable cross-chain audit workflow and prioritization queues for security operations.